# Explaining Sources of Uncertainty in Automated Fact-Checking

### v1

@@ Acknowledgments
[Uncaptioned image][Uncaptioned $\end{array}$ This research was co-funded by the European Union (ERC, ExplainYourself, 101077481), by the Pioneer Centre for AI, DNRF grant number P1, as well as by The Villum Synergy Programme. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Backbone model performance on public benchmarks
OLMo-2-1124-13B-Instruct (Team OLMo et al., 2024) 13 B 67.5 54.2
@@ Appendix B Method: Selecting attention heads to steer
Appendix C Prompt Example for Assigning Relation Labels to Captured Span Interactions
@@ 3. Output the span pairs exactly as given, followed by
Appendix D Perturbation details for faithfulness measurement
@@
To evaluate how faithfully each NLE reflects model uncertainty, we generate multiple counterfactuals per instance, following Atanasova et al. (2020) and Siegel et al. (2024) (see §5.1). For every input, comprising one claim and two evidence passages, we first tag part-of-speech with spaCy, then choose seven random insertion sites. At each site we insert either (i) a random adjective before a noun or (ii) a random adverb before a verb. The candidate modifiers are drawn uniformly from the full WordNet lists of adjectives and adverbs. Because we sample three random candidates for each of the four positions, this procedure yields $4\times 3=12$ perturbations per instance, providing a sufficient set for the subsequent Entropy-CCT evaluation, in which we check whether the NLE mentions the inserted word and correlate that mention with the uncertainty change induced by each perturbation.
Appendix E Differences Between Entropy-CCT and CCT
In CCT test, Total Variation Distance (TVD) is computed between two probability distributions $P$ and $Q$ as $\text{TVD}(P,Q)=\frac{1}{2}\sum_{i}|P_{i}-Q_{i}|$, measuring the absolute change in class-wise probabilities. We instead operate on the entropies of those distributions, yielding a single-valued measure of uncertainty shift.
@@ Appendix E Differences Between Entropy-CCT and CCT
Appendix F Prompt template for PromptBaseline, CLUE-Span and CLUE-Span+Steering on Healthver and Druid dataset
@@
F.1 Prompt template for PromptBaseline
@@ 2. Explain your predictions uncertainty by identifying the three most
Figure 4: Three-shot prompt for PromptBaseline (Shots 2–3 omitted) on the HealthVer and DRuiD datasets.
F.2 Prompt template for CLUE-Span and CLUE-Span+Steering
@@ 2. Explain your predictions uncertainty by referring to the three span
Figure 5: Three-shot prompt for CLUE-Span and CLUE-Span+Steering (Shots 2–3 omitted) on the HealthVer and Druid datasets.
Appendix G Extended Statistical Analysis of Faithfulness Scores
This section elaborates on the statistical evaluation of faithfulness regarding (i) recalling the definition and intuitive interpretation of the point–biserial coefficient $r_{\text{pb}}$(E.q. 9), (ii) outlining the $t$-test used to assess significance, (iii) reporting the faithfulness results (§5.1) along with statistical results. Note that, each dataset is evaluated on $n=600\times 12=7{,}200$ perturbations with 600 instances with 12 perturbations each (see App. D). and (iv) demonstrating through concise numerical summaries that both CLUE-Span and CLUE-Span+Steering are significantly more faithful than the PromptBaseline.
G.1 Interpreting $r_{\text{pb}}$ and $\Delta r_{\text{pb}}$
The Entropy-CCT score is the point-biserial correlation (Tate, 1954) between the absolute entropy change $|\Delta u|$ and the binary mention flag $m$. Because it is mathematically identical to a Pearson $r$ computed between one continuous and one binary variable, it obeys $-1\leq r_{\text{pb}}\leq 1$. When $r_{\text{pb}}=0$, it means the high- and low-impact perturbations are mentioned equally often. If the two strata are roughly balanced, every $+0.01$ in $r_{\text{pb}}$ increases the probability that a truly uncertainty-influential token is mentioned by about one percentage point (pp). A gain $\Delta r_{\text{pb}}$ therefore translates to an absolute improvement of $\approx|\Delta r_{\text{pb}}|\times 100$,pp in mention rate. For instance, moving from $-0.08$ to $+0.06$ is a swing of $0.14$, corresponding to, $14$,pp.
G.2 Significance testing
Because the point-biserial is a Pearson correlation, the familiar $t$–test applies:
$t$ $=r_{\text{pb}}\,\sqrt{\frac{n-2}{1-r_{\text{pb}}^{2}}},$ (12)
$t$ $\sim t_{(n-2)}\qquad\text{under }H_{0}\!:r_{\text{pb}}=0.$ (13)
G.3 Faithfulness with significance results
Table  4 shows the point-biserial coefficients $r_{\text{pb}}$, which is our faithfulness measurement for model uncertainty(See, E.q.9), the associated $t$ statistics, and two-sided $p$ values for every model–method pair. Values that meet the stricter $p<0.01$ criterion are highlighted in bold.
Model Method $r_{\text{pb}}$ $t$ $p$
HealthVer
Qwen2.5-14B-Instruct PromptBaseline $-0.028$ $-2.38$ $1.7\times 10^{-2}$
CLUE-Span $+0.006$ $+0.51$ $6.1\times 10^{-1}$
CLUE-Span+Steering $+0.033$ $+2.80$ $\mathbf{5.1\times 10^{-3}}$
OLMo-2-1124-13B-Instruct PromptBaseline $-0.100$ $-8.53$ $\mathbf{<10^{-15}}$
CLUE-Span $+0.005$ $+0.42$ $6.7\times 10^{-1}$
CLUE-Span+Steering $+0.020$ $+1.70$ $9.0\times 10^{-2}$
Gemma-2-9B-IT PromptBaseline $-0.105$ $-8.96$ $\mathbf{<10^{-15}}$
CLUE-Span $+0.007$ $+0.59$ $5.5\times 10^{-1}$
CLUE-Span+Steering $+0.021$ $+1.78$ $7.5\times 10^{-2}$
DRUID
Qwen2.5-14B-Instruct PromptBaseline $-0.080$ $-6.81$ $\mathbf{9.8\times 10^{-12}}$
CLUE-Span $+0.089$ $+7.58$ $\mathbf{3.4\times 10^{-14}}$
CLUE-Span+Steering $+0.102$ $+8.70$ $\mathbf{<10^{-15}}$
OLMo-2-1124-13B-Instruct PromptBaseline $-0.130$ $-11.12$ $\mathbf{<10^{-15}}$
CLUE-Span $+0.014$ $+1.19$ $2.3\times 10^{-1}$
CLUE-Span+Steering $+0.099$ $+8.44$ $\mathbf{<10^{-15}}$
Gemma-2-9B-IT PromptBaseline $-0.120$ $-10.26$ $\mathbf{<10^{-15}}$
CLUE-Span $+0.043$ $+3.65$ $\mathbf{2.6\times 10^{-4}}$
CLUE-Span+Steering $+0.098$ $+8.35$ $\mathbf{<10^{-15}}$
Table 4: Detailed faithfulness evaluation results for baseline method PromptBaseline, and two variants of our CLUE framework CLUE-Span and CLUE-Span+Steering on Healthver and Druid dataset based on Qwen2.5-14B-Instruct(Qwen Team (2024)), OLMo-2-1124-13B-Instruct(Team OLMo et al. (2024))and Gemma-2-9B-IT(Gemma Team (2024)). Point-biserial correlation $r_{\text{pb}}$ is our Entropy-CCT measurement(§5.1), along with $t$ statistic and two-sided $p$-value for each model–method pair ($n=7{,}200$, $df=7{,}198$). Entries with $p<0.01$ are bold.
Across both datasets and all three backbones, the PromptBaseline exhibits negative correlations, implying an non-faithful tendency to highlight low-impact tokens within the generation NLEs, with $\text{mean}=-0.094$. The prompt-only variant of our CLUE framework CLUE-Span neutralises this bias and turns the average into $+0.027$; three of its six coefficients are clear $p<0.01$, indicating a modest but significant improvement regarding faithfulness.
The full CLUE-Span+Steering variant pushes the mean to $+0.062$ and achieves $p<0.01$ in four of six settings. Interpreting these numbers via  §G.1, the switch from $-0.094$ to $+0.062$ yields a absolute increase of (0.062(0.094))×100!!16(0.062-(-0.094)$\times$ 100$\approx$!16( 0.062 - ( - 0.094 ) ) × 100 ! ≈ ! 16, pp in the probability that a truly influential token of uncertainty is named in the NLE, which is easily noticeable in qualitative inspection.
Appendix H Human Evaluation Details
H.1 Participants and Materials

### v2

@@ Acknowledgments
[…]
@@ Appendix A Backbone model performance on public benchmarks
OLMo-2-1124-13B-Instruct (Team OLMo et al., 2025) 13 B 67.5 54.2
@@ Appendix B Method: Selecting attention heads to steer
Appendix C Conflict and Agreement Identification Details
C.1 Prompt Example for Assigning Relation Labels to Captured Span Interactions
@@ 3. Output the span pairs exactly as given, followed by
C.2 Open-weight Alternative for Span-Interaction Relation Labeling
This section studies how the choice of relation labeler affects downstream NLE quality in CLUE. While we use GPT-4o as the default relation labeler to reduce labeling noise, we show that a strong open-weight alternative is also feasible, with only a small performance drop.
As discussed in Sec. 2.3, CLUE uses a relation labeler $L_{\mathrm{rel}}$ to assign each extracted span interaction one of $\{\textsc{agree},\textsc{disagree},\textsc{unrelated}\}$. We considered two labeler options: (i) a closed-API model, which can reduce label noise at the cost of API usage, and (ii) an open-weight model, which is fully local but may yield slightly noisier labels and thus introduce additional noise in downstream NLE generation.
To quantify the impact of the labeler choice, we reran the pipeline with one open-weight alternative for $L_{\mathrm{rel}}$: Qwen2.5-14B-Instruct, keeping all other components fixed. Table 4 reports the results.
HealthVer DRUID
Relation labeler Faith. ($\uparrow$) Span-Cov. ($\uparrow$) Span-Ext. ($\downarrow$) LEE ($\uparrow$) Faith. ($\uparrow$) Span-Cov. ($\uparrow$) Span-Ext. ($\downarrow$) LEE ($\uparrow$)
GPT-4o 0.033 0.44 0.53 0.80 0.102 0.28 0.20 0.77
Qwen2.5-14B-Instruct 0.025 0.41 0.58 0.78 0.076 0.25 0.25 0.75
Table 4: Effect of span-interaction relation labeler choice. We rerun CLUE while swapping only $L_{\mathrm{rel}}$ (GPT-4o vs. Qwen2.5-14B-Instruct) and keeping all other components fixed (backbone: Qwen2.5-14B-Instruct; variant: CLUE-Span+Steering; k=3k{=}3). An open-weight labeler is feasible with only a small drop in downstream NLE quality. Bold indicates the better result per metric within each dataset.
Overall, Qwen2.5-14B-Instruct yields similar, though slightly weaker, downstream NLE quality compared to GPT-4o across both datasets. For instance, on HealthVer, Faithfulness drops from $0.033$ to $0.025$ and LEE from $0.80$ to $0.78$; on DRUID, Faithfulness drops from $0.102$ to $0.076$ and LEE from $0.77$ to $0.75$.
These results indicate that an open-weight labeler is feasible in CLUE, and the framework is not tied to a proprietary component. In this work, we use GPT-4o in the main experiments because it provides consistently stronger relation labeling and reduces the risk of error propagation from noisy relation labels to the final explanations. We further discuss how upstream labeling errors can propagate and mitigation strategies in App. C.3.
A natural concern is whether using a closed-API labeler compromises transparency. We emphasize that CLUE is white-box with respect to the backbone model: uncertainty scoring, span-interaction extraction, and attention steering require access to the backbone model internals, which closed APIs do not provide. Accordingly, running the entire pipeline with GPT-4o would forgo these white-box signals and reduce the method to surface-level prompting; prior work shows such verbalized uncertainty can be unreliable, and fact-checking practice requires systems to “show the work” behind a decision (Yona et al., 2024; Tanneru et al., 2024; Warren et al., 2025).
GPT-4o is used only for the narrow task of pairwise relation judgment after span interactions have already been extracted from backbone internals; as such, it does not affect the transparency of the core framework.
C.3 Error Propagation Analysis
We manually inspected 50 HealthVer instances (top-$3$ span pairs per instance; N=150N{=}150 span pairs) to assess (i) span extraction quality and (ii) relation labeling accuracy when using GPT-4o as $L_{\mathrm{rel}}$.
Span extraction quality.
Most extracted spans are very short: 105/150 span pairs contain two-token spans (often salient nouns such as subjects/objects). In 35/150 cases, at least one span is incomplete (e.g., “D” instead of “vitamin D”) or semantically underspecified (e.g., “is not”). Overall, the extractor tends to favor short, high-salience tokens, suggesting potential for future work to encourage more complete semantic units.
Relation labelling accuracy.
On these 150 span pairs, GPT-4o achieves 86% accuracy (129/150 correct). Among the 21 errors, 12 are cases where an agree relation is mislabelled as unrelated, typically when one span is incomplete (e.g., “D” vs. “vitamin D”), indicating that span completeness is a primary failure mode.
Propagation to downstream NLEs.
Using Qwen2.5-14B-Instruct with CLUE-Span+Steering, 13/21 mislabeled interactions are explicitly mentioned in the generated NLEs, showing that upstream errors can propagate into explanations. Moreover, the subset of instances with mislabeled interactions exhibits lower faithfulness (mean Entropy-CCT 0.020) than the overall HealthVer–Qwen CLUE-Span+Steering average (0.033; Table 1), suggesting that incorrect relation labels can reduce explanation faithfulness.
This analysis suggests that fragmentary spans and relation-label noise can be an upstream error source and carry forward to the final NLEs, motivating mitigation strategies discussed in App. C.4 for future work.
C.4 Mitigating Error Propagation
We outline several directions to mitigate error propagation in future work: (1) Improve span extraction: add mild constraints during community detection so extracted spans more consistently capture complete semantic units rather than clipped tokens. (2) Improve relation labeling robustness: use stronger prompting (broader local context, higher-quality example shots) or lightly fine-tune an open-weight model on a compact relation-judgment set. (3) Generation-time sanity checks: add a lightweight consistency check that flags or down-weights span pairs judged internally inconsistent with the claim/evidence before producing the final NLE.
Appendix D Resources and Model Cards
Appendix E Computational Cost and Resource Requirements
Setup.
We report the computational overhead of CLUE compared to PromptBaseline. Unless otherwise stated, measurements are collected using Qwen2.5-14B-Instruct on the HealthVer dataset with a single NVIDIA A100 80GB GPU. All values are reported per instance (averaged over the evaluation set).
E.1 Latency
Table 5 reports the average wall-clock time per instance for the three methods. Both CLUE-Span and CLUE-Span+Steering are roughly 2 seconds slower per instance than PromptBaseline, and this overhead is entirely due to the additional span-interaction extraction and relation labeling stage.
Approach Uncertainty Scoring (s) Span Extraction + Labeling (s) NLE Generation (s) Total (s)
PromptBaseline 0.364 14.8969 15.2609
CLUE-Span 0.364 2.0452 14.7730 17.1822
CLUE-Span+Steering 0.364 2.0452 14.7874 17.1966
Table 5: Average time cost per instance for the three NLE generation methods on HealthVer using Qwen2.5-14B-Instruct on a single NVIDIA A100 80GB GPU. “–” indicates that span extraction/labeling is not used by PromptBaseline.
E.2 GPU memory usage
Table 6 reports peak GPU memory usage (VRAM). During NLE generation, VRAM usage is similar across all methods. CLUE-Span+Steering requires about 0.32 GiB more VRAM than PromptBaseline, primarily because the prompt is longer (it includes the extracted span interactions; see App. G.2). Notably, CLUE-Span+Steering does not require more VRAM than CLUE-Span, since during generation KV-cache and activations dominate memory and steering does not add a separate memory footprint beyond the base forward pass.
Approach Uncertainty Scoring (GiB) Span Interaction Extraction (GiB) NLE Generation (GiB)
PromptBaseline 29.71 30.52
CLUE-Span 29.71 29.72 30.84
CLUE-Span+Steering 29.71 29.72 30.84
Table 6: Average peak VRAM per instance for the three NLE generation methods (HealthVer, Qwen2.5-14B-Instruct, NVIDIA A100 80GB).
E.3 Memory requirement for steering 100 heads
For selecting the most important 100 attention heads (App. B), the peak VRAM is 29.8 GiB. For steering these heads during NLE generation (CLUE-Span+Steering), the average peak VRAM per instance is 30.84 GiB. Thus, a GPU with 40GB of memory satisfies the requirement for attention steering across 100 heads.
E.4 GPT-4o API cost for relation labeling
When using GPT-4o as $L_{\text{rel}}$, the total relation-labeling cost scales linearly with the number of labeled instances. In our main experiments, the cumulative cost across datasets and backbones was approximately \$30 (about \$0.003 per instance), with an average input length of $\sim$1k tokens and an average output length of $\sim$50 tokens per call.
Appendix F Methodology Supplement: Faithfulness Evaluation Details
This supplement provides implementation and statistical details for the Entropy-CCT faithfulness metric introduced in §4.1. We describe (i) how counterfactual perturbations are generated, (ii) how Entropy-CCT relates to the original CCT, and (iii) how we test statistical significance and report results.
F.1 Entropy-CCT Perturbation Generation
@@
To evaluate how faithfully each NLE reflects model uncertainty, we generate multiple counterfactuals per instance, following Atanasova et al. (2020) and Siegel et al. (2024) (see §4.1). For every input, comprising one claim and two evidence passages, we first tag part‐of‐speech with spaCy (Honnibal and Montani, 2017) (en_core_web_sm), then choose four random insertion sites (uniformly among tokens tagged as NOUN or VERB). At each site we insert either (i) a random adjective before a noun or (ii) a random adverb before a verb. The candidate modifiers are drawn uniformly from WordNet (Miller, 1992) using NLTK (Bird et al., 2009): we construct adjective/adverb candidate lists by collecting all lemma names from all WordNet adjective/adverb synsets, respectively. Because we sample three random candidates for each of the four positions, this procedure yields $4\times 3=12$ perturbations per instance, providing a sufficient set for the subsequent Entropy-CCT evaluation, in which we check whether the NLE mentions the inserted word and compute the correlation between that mention and the uncertainty change induced by each perturbation.
F.2 Entropy‑CCT vs. Original CCT
CCT Siegel et al. (2024) evaluates faithfulness by perturbing the input and measuring the resulting change in the model’s predicted label distribution. Specifically, it uses the Total Variation Distance (TVD) between the pre- and post-perturbation distributions $P(\cdot\mid X)$ and $P(\cdot\mid X^{\prime})$:
$\mathrm{TVD}(P,Q)=\frac{1}{2}\sum_{i}|P_{i}-Q_{i}|.$ (12)
Entropy-CCT is a minimal adaptation of CCT that preserves the same perturbation procedure and the same correlational test: we correlate the perturbation impact with whether the inserted token is mentioned in the NLE. The only change is the signal we correlate. Instead of TVD (probability-shift), we use Absolute Entropy Change (AEC), i.e., the change in predictive entropy, to directly target uncertainty (our object of explanation).
F.3 Statistical Testing & Significance
We test statistical significance of $r_{\text{pb}}$ using the standard Pearson correlation $t$-test. Each perturbation constitutes one observation: for each dataset we evaluate 600 instances with 12 perturbations per instance, yielding $n=600\times 12=7{,}200$ observations (App. F.1). Because the point-biserial correlation is algebraically equivalent to Pearson correlation between a continuous and a binary variable, we use the standard Pearson $t$-test for $H_{0}:r_{\text{pb}}=0$.
Interpreting $r_{\text{pb}}$ and $\Delta r_{\text{pb}}$. The Entropy-CCT score is the point-biserial correlation (Tate, 1954) between the absolute entropy change $|\Delta u|$ and the binary mention flag $m$. Because it is mathematically identical to a Pearson $r$ computed between one continuous and one binary variable, it obeys $1\leq r_{\text{pb}}\leq 1$. When $r_{\text{pb}}=0$, it means the high- and low-impact perturbations are mentioned equally often. If the two strata are roughly balanced, every +0.01+0.01 in $r_{\text{pb}}$ increases the probability that a truly uncertainty-influential token is mentioned by about one percentage point (pp). A gain $\Delta r_{\text{pb}}$ therefore translates to an absolute improvement of $\approx|\Delta r_{\text{pb}}|\times 100$ pp in mention rate. For instance, moving from 0.08-0.08 to +0.06+0.06 is a swing of $0.14$, corresponding to about $14$ pp.
Significance testing. Because the point-biserial is a Pearson correlation, the familiar tt–test applies:
$t$ $=r_{\text{pb}}\,\sqrt{\frac{n-2}{1-r_{\text{pb}}^{2}}},$ (13)
$t$ $\sim t_{(n-2)}\qquad\text{under }H_{0}\!:r_{\text{pb}}=0.$ (14)
Faithfulness with significance results. Table  7 shows the point-biserial coefficients $r_{\text{pb}}$, which is our faithfulness measurement for model uncertainty(See, Eq.9), the associated $t$ statistics, and two-sided $p$ values for every model–method pair. Values that meet the stricter $p<0.01$ criterion are highlighted in bold.
Model Method $r_{\text{pb}}$ $t$ $p$
HealthVer
Qwen2.5-14B-Instruct PromptBaseline 0.028-0.028 2.38-2.38 $1.7\times 10^{-2}$
CLUE-Span +0.006+0.006 +0.51+0.51 $6.1\times 10^{-1}$
CLUE-Span+Steering +0.033+0.033 +2.80+2.80 $\mathbf{5.1\times 10^{-3}}$
OLMo-2-1124-13B-Instruct PromptBaseline 0.100-0.100 8.53-8.53 $\mathbf{<10^{-15}}$
CLUE-Span +0.005+0.005 +0.42+0.42 $6.7\times 10^{-1}$
CLUE-Span+Steering +0.020+0.020 +1.70+1.70 $9.0\times 10^{-2}$
Gemma-2-9B-IT PromptBaseline 0.105-0.105 8.96-8.96 $\mathbf{<10^{-15}}$
CLUE-Span +0.007+0.007 +0.59+0.59 $5.5\times 10^{-1}$
CLUE-Span+Steering +0.021+0.021 +1.78+1.78 $7.5\times 10^{-2}$
DRUID
Qwen2.5-14B-Instruct PromptBaseline 0.080-0.080 6.81-6.81 $\mathbf{9.8\times 10^{-12}}$
CLUE-Span +0.089+0.089 +7.58+7.58 $\mathbf{3.4\times 10^{-14}}$
CLUE-Span+Steering +0.102+0.102 +8.70+8.70 $\mathbf{<10^{-15}}$
OLMo-2-1124-13B-Instruct PromptBaseline 0.130-0.130 11.12-11.12 $\mathbf{<10^{-15}}$
CLUE-Span +0.014+0.014 +1.19+1.19 $2.3\times 10^{-1}$
CLUE-Span+Steering +0.099+0.099 +8.44+8.44 $\mathbf{<10^{-15}}$
Gemma-2-9B-IT PromptBaseline 0.120-0.120 10.26-10.26 $\mathbf{<10^{-15}}$
CLUE-Span +0.043+0.043 +3.65+3.65 $\mathbf{2.6\times 10^{-4}}$
CLUE-Span+Steering +0.098+0.098 +8.35+8.35 $\mathbf{<10^{-15}}$
Table 7: Detailed faithfulness evaluation results for baseline method PromptBaseline, and two variants of our CLUE framework CLUE-Span and CLUE-Span+Steering on Healthver and Druid dataset based on Qwen2.5-14B-Instruct(Qwen Team (2024)), OLMo-2-1124-13B-Instruct(Team OLMo et al. (2025))and Gemma-2-9B-IT(Gemma Team (2024)). Point-biserial correlation $r_{\text{pb}}$ is our Entropy-CCT measurement(§4.1), along with $t$ statistic and two-sided $p$-value for each model–method pair (n=7,200n=7{,}200, df=7,198df=7{,}198). Entries with $p<0.01$ are bold.
Across both datasets and all three backbones, the PromptBaseline exhibits negative correlations, implying a non-faithful tendency to highlight low-impact tokens in the generated NLEs, with $\text{mean}=-0.094$. The prompt-only variant of our CLUE framework CLUE-Span neutralises this bias and turns the average into +0.027+0.027; two of its coefficients are clear $p<0.01$, indicating a modest but significant improvement regarding faithfulness.
The full CLUE-Span+Steering variant pushes the mean to +0.062+0.062 and achieves $p<0.01$ in four of six settings. Interpreting these numbers using the discussion above, the switch from 0.094-0.094 to +0.062+0.062 yields an absolute increase of $(0.062-(-0.094))\times 100!\approx 16$, pp in the probability that a truly uncertainty-influential token is named in the NLE, which is easily noticeable in qualitative inspection.
@@ F.3 Statistical Testing & Significance
Appendix G Prompt template for PromptBaseline, CLUE-Span and CLUE-Span+Steering on Healthver and Druid dataset
@@
G.1 Prompt template for PromptBaseline
@@ 2. Explain your predictions uncertainty by identifying the three most
Figure 4: Three-shot prompt for PromptBaseline (Shots 2-3 omitted) on the HealthVer and DRUID datasets.
G.2 Prompt template for CLUE-Span and CLUE-Span+Steering
@@ 2. Explain your predictions uncertainty by referring to the three span
Figure 5: Three-shot prompt for CLUE-Span and CLUE-Span+Steering (Shots 2-3 omitted) on the HealthVer and DRUID datasets.
Appendix H Extended Evaluation Results
H.1 Extended Evaluation Results for Scenarios Involving Three Evidence pieces
Metrics PromptBase CLUE-S CLUE-Steer
Faith. ($\uparrow$) 0.08 0.10 0.13
Span-Cov. ($\uparrow$) 0.25 0.41
Span-Ext. ($\downarrow$) 0.23 0.27
LEE ($\uparrow$) 0.80 0.86 0.86
Table 8: Automatic evaluation for three types of NLE on HealthVer using Qwen2.5-14B-Instruct when one claim and three pieces of evidence are involved: PromptBase(PromptBaseline) is the baseline method, CLUE-S (CLUE-Span) and CLUE-Steer(CLUE-Span+Steering) are the two variants of our CLUE framework.
To verify the generalizability of our framework in scenarios involving multiple evidence pieces, we conduct an experiment on the Healthver dataset using the Qwen2.5‑14B‑Instruct model, pairing each claim with three pieces of evidence instead of two (as shown in Tab. 8). Note that we select the top‑$k=6$ span interactions for NLE generation, taking the top‑1 interaction from each pairwise combination among the $C$, $E_{1}$, $E_{2}$, and $E_{3}$ input parts.
Overall, we observe a similar trend with three pieces of evidence as with two (see §4.5): CLUE-Span+Steering and CLUE-Span show better Faithfulness and Label-Explanation Entailment than PromptBaseline. Interestingly, the Faithfulness of all three explanation types increases relative to the two‑evidence setting, from -0.028 to 0.08 for PromptBaseline, from 0.006 to 0.10 for CLUE-Span, and from 0.033 to 0.13 for CLUE-Span+Steering. This may be because, when NLEs are grounded in richer evidence, they capture more nuances in both claim-evidence and inter‑evidence interactions, thereby more faithfully reflecting the model’s fact‑checking decision.
It is notable that CLUE-Span+Steering includes more targeted span interactions (higher Span‑Coverage: 0.41 > 0.27) as well as more redundant span interactions (higher Span‑Extraneous: 0.27 > 0.23) than CLUE-Span. This suggests that, although CLUE-Span+Steering better focuses on target interactions, it still suffers from redundancy due to the inclusion of more irrelevant interactions, which we call for future work to address.
H.2 Extended Evaluation Results on Reasoning Model
To demonstrate that our framework generalizes to reasoning models trained primarily via large‑scale reinforcement learning, and designed to encourage chain‑of‑thought exploration, we adopt DeepSeek‑R1‑Distill‑Qwen‑14B [^5] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-14B (DeepSeek-AI, ) as the base model and evaluate it on the HealthVer dataset in the setting with one claim and two pieces of evidence (see Tab. 9).
Metrics PromptBase CLUE-S CLUE-Steer
Faith. ($\uparrow$) 0.12 0.16 0.18
Span-Cov. ($\uparrow$) 0.58 0.67
Span-Ext. ($\downarrow$) 0.59 0.64
LEE ($\uparrow$) 0.88 0.90 0.89
Table 9: Automatic evaluation for three types of NLE on HealthVer using DeepSeek-R1-Distill-Qwen-14B when one claim and two pieces of evidence are involved: PromptBase(PromptBaseline) is the baseline method, CLUE-S (CLUE-Span) and CLUE-Steer(CLUE-Span+Steering) are the two variants of our CLUE framework.
Overall, both variants of our CLUE model, CLUE-Span and CLUE-Span+Steering, show higher Faithfulness and Label‑Explanation Entailment than the baseline method PromptBaseline, confirming the effectiveness of our framework. It is also notable that, compared with the evaluation results on other models in Tab. 1, all three types of explanations generated with the DeepSeek‑R1‑Distill‑Qwen‑14B model achieve the best performance across metrics, especially the highest Faithfulness and Span‑Coverage. The former implies that reinforcement-learning-trained reasoning models may be more capable of generating highly faithful explanations with the thinking step; the latter suggests that this type of model follows our interaction-guided instructions more effectively. Consequently, both CLUE-Span and CLUE-Span+Steering include more target-span interactions, yielding higher Span-Coverage.
Comparing the two variants, CLUE-Span+Steering exhibits higher Span‑Coverage (0.67 > 0.58) than CLUE-Span, but also more Span‑Extraneous (0.64 > 0.59), suggesting that CLUE-Span+Steering tends to include redundant interactions while focusing on target interactions. CLUE-Span+Steering also achieves the highest Faithfulness (0.18 > 0.16), consistent with its higher Span‑Coverage, but yields similar Label‑Explanation Entailment to CLUE-Span (0.89 vs. 0.90), likely due to the increased Span‑Extraneous. Addressing this trade‑off is left for future work.
Appendix I Human Evaluation Details
I.1 Participants and Materials